OpenStudy (anonymous):

2/3 + 3/8 = 5/8 5/24 11/24 Not Here 1 1/24

OpenStudy (hoblos):

2/3 + 3/8 = 16/24 + 9/24 = 25/24 = 1 1/24

OpenStudy (anonymous):

thank you !!!

OpenStudy (hoblos):

my pleasure XD

Directrix (directrix):

1 1/24 2/3 + 3/8 = 16/24 + 9/24 = 25/24 = 1 and 1/24